### v1

# Scene123: One Prompt to 3D Scene Generation via Video-Assisted and Consistency-Enhanced MAE

Abstract
As Artificial Intelligence Generated Content (AIGC) advances, a variety of methods have been developed to generate text, images, videos, and 3D objects from single or multimodal inputs, contributing efforts to emulate human-like cognitive content creation. However, generating realistic large-scale scenes from a single input presents a challenge due to the complexities involved in ensuring consistency across extrapolated views generated by models. Benefiting from recent video generation models and implicit neural representations, we propose Scene123, a 3D scene generation model, that not only ensures realism and diversity through the video generation framework but also uses implicit neural fields combined with Masked Autoencoders (MAE) to effectively ensures the consistency of unseen areas across views. Specifically, we initially warp the input image (or an image generated from text) to simulate adjacent views, filling the invisible areas with the MAE model. However, these filled images usually fail to maintain view consistency, thus we utilize the produced views to optimize a neural radiance field, enhancing geometric consistency. Moreover, to further enhance the details and texture fidelity of generated views, we employ a GAN-based Loss against images derived from the input image through the video generation model. Extensive experiments demonstrate that our method can generate realistic and consistent scenes from a single prompt. Both qualitative and quantitative results indicate that our approach surpasses existing state-of-the-art methods. We show encourage video examples at https://yiyingyang12.github.io/Scene123.github.io/.
Figure 1: Some examples generated by Scene123. For a single input image or text, our method can generate 3D scenes with consistent perspectives, fine geometry, and realistic textures, applicable to real, virtual, or object-centered scenes.
1 Introduction
3D scene generation aims to create realistic or stylistically specific scenes from limited prompts, such as a few images or a text description. This is a fundamental issue in computer vision and graphics and a critical challenge in generative artificial intelligence. Recent advancements have demonstrated substantial progress through the use of vision-language models [58, 37], generative models such as Generative Adversarial Networks (GANs) [96], Variational autoencoders (VAEs) [67], or diffusion models [88], and scene representations like Neural Radiance Fields (NeRF) [49] or 3D Gaussian Splatting (3DGS) [36]. Typically, these methods begin by generating images with a pretrained generative model [63, 90], or by directly using images for image-to-scene generation, then estimate additional 3D surface geometric details such as depth and normal [54, 38], and subsequently render the surface textures of the scenes using generative strategies like inpainting [85, 1]. Moreover, some approaches update the geometric surfaces as new views are generated to maintain the coherence of the scene [95]. However, these methods often rely on pre-trained models [58, 37], resulting in inconsistencies and artifacts in the generated scenes. Additionally, these methods face challenges in producing high-quality, coherent 3D representations across diverse and complex environments.
In this paper, we endeavor to address this challenge about generating 3D scenes from a single image or textual description, ensuring viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, as shown in Fig. 1. However, achieving a balance between viewpoint consistency and flexibility presents a significant challenge. View consistency requires maintaining coherent and accurate details across multiple perspectives, which can limit the model’s adaptability to diverse inputs and tasks. Conversely, flexibility requires the model to produce high-quality outputs under varying conditions, which may introduce inconsistencies in the generated scenes.
It is noteworthy that scene synthesis based on multi-view images is a long-studied topic from the early Multi-View Stereo (MVS)  [70, 69] to recent implicit neural representations [92, 2], ensuring view consistency in generated scenes through multi-view matching mechanisms. However, when the input is reduced to a single image or a textual description, lacking reference multi-view images, the performance of these methods is greatly compromised. Fortunately, methodologies such as Masked Autoencoders (MAE)  [23] provide avenues for extrapolating to areas unseen in new views, while the incorporation of additional semantic layers augments the coherence of the synthesized scenes. Additionally, video generation models  [26, 44] facilitate the enhancement of scenes with richer priors and more detailed texture information. Thus, how to utilize multi-view reconstruction with robust physical constraints alongside effective expansions in new viewpoints and video generation models is still an unreached area.
Scene123 investigates a methodology for 3D scene reconstruction employing stringent physical constraints alongside a robust multi-view MAE and video generation models to ensure view consistency. To achieve this, for each input or generated single image, we first create images of nearby perspectives through warping. Subsequently, a consistency-enhanced MAE is designed to inpaint the unseen areas, with a shared codebook maintained to distribute global information to every invisible area. We implement a progressive strategy derived from Text2NeRF  [95] to incrementally update perspectives throughout this process. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images. The synergistic operation of these two modules enables Scene123 to generate consistent, finely detailed three-dimensional scenes with photo-realistic textures from a single prompt.
We conduct extensive experiments on text-to-scene and image-to-scene generation, encompassing both real and virtual scenes. The results offer robust empirical evidence that strongly supports the effectiveness of our framework. Our contributions can be summarized as: 1) We propose a novel scene generation framework based on one prompt, which establishes the connection between MAE and video generation models for the first time to ensure view consistency and realism of the generated scenes. 2) The Consistency-Enhanced MAE is designed to fill unseen areas in new views by injecting global semantic information and combining it with neural implicit fields, ensuring consistent surface representation across various views. 3) We introduce the video-assisted 3D-aware generative refinement module, which enhances scene reconstruction by integrating the diversity and realism of video generation models through a GAN-based function to significantly improve detail and texture fidelity. 4) Extensive experiments validate the efficacy of Scene123, demonstrating greater accuracy in surface reconstruction, higher realism in reconstructed views, and better texture fidelity compared to the state-of-the-art methods. The code and weights will be available.
2 Related Work
3D Scene Generation. Recent advancements in text-to-3D and image-to-3D scene generation have been remarkable. Most research in this domain can be classified into two main categories: object-centric [14, 34, 40, 41, 47, 56, 57, 60, 72, 74, 78, 86, 92], which focuses on objects without backgrounds, and holistic, which involves generating entire 3D scenes or 3D trajectories including backgrounds. Expanding these techniques to create fully unconstrained 3D scenes remains a significant challenge. Many recent advancements in text-driven 3D scene generation have focused on modeling 3D scenes using text inputs [32, 13, 95, 19, 28, 94, 73, 97]. Due to the scarcity of paired text-3D scene data, most studies utilize Contrastive Language-Image Pre-training (CLIP) [58] or pre-trained text-to-image models to interpret the text input. Text2Scene [32] employs CLIP to model and stylize 3D scenes from text (or image) inputs by decomposing the scene into manageable sub-parts. Set-the-Scene [13] and Text2NeRF [95] generate NeRFs from text using text-to-image diffusion models to represent 3D scenes. SceneScape [19] and Text2Room [28] leverage a pre-trained monocular depth prediction model for enhanced geometric consistency and directly generate the 3D textured mesh representation of the scene. However, these methods depend on inpainted images for scene completion, which, while producing realistic visuals, suffer from limited 3D consistency. More recent studies [20, 2, 80] have successfully generated multi-object compositional 3D scenes. Another class of methods utilizes auxiliary inputs, such as layouts [55, 17], to enhance scene generation. Unlike text captions, which can be rather vague, some approaches generate 3D scenes from image inputs, where the 3D scene corresponds closely to the depicted image. Early scene generation methods often require specific scene data for training to obtain category-specific scene generators, such as GAUDI [3], or focus on single scene reconstruction based on the input image, such as PixelSynth [61] and Worldsheet [31]. However, these methods are often limited by the quality of the generation or the extensibility of the scene.
Recently, numerous studies have concentrated on generating 3D scenes from image inputs. PERF [83] generates 3D scenes from a single panoramic image, using diffusion models to supplement shadow areas. ZeroNVS [68] extends this capability by reconstructing both objects and environments in 3D from a single image. Despite its environmental reconstruction lacking some detail, the algorithm demonstrates an understanding of environmental contexts. LucidDreamer [12] and WonderJourney [93] utilize general-purpose depth estimation models to project hallucinated 2D scene extensions into 3D representations. However, these methods still face challenges in achieving realism, often producing artifacts due to the reliance on pre-trained models.
Video Diffusion Models and 3D-aware GANs. Diffusion models [76] have recently emerged as powerful generative models capable of producing a wide array of images [5, 63, 65] and videos [4, 22, 81] by iteratively denoising a noise sample. Among these models, the publicly available Stable Diffusion (SD) [63] and Stable Video Diffusion (SVD) [4] exhibit strong generalization capabilities due to training on extremely large datasets such as LAION [71] and LVD [4]. Consequently, they are frequently used as foundational models for various generation tasks, including novel view synthesis. To enhance generalization and multi-view consistency, some contemporary works leverage the temporal priors in video diffusion models for object-centric 3D generation. For instance, IM-3D [46] and SV3D [82] explore the capabilities of video diffusion models in object-centric multi-view generation. V3D [11] extends this approach to scene-level novel view synthesis. However, these methods often produce unsatisfactory results for complex objects or scenes, leading to inconsistencies among multiple views or unrealistic geometries.
Early works, like 3D-GAN [87], Pointflow [89], and ShapeRF [6] focus more on the category-specific texture-less geometric shape generation based on the representations of voxels or point clouds. However, limited by the generation capabilities of GANs, these methods can only generate rough 3D assets of specific categories. Subsequently, HoloGAN [51], GET3D [21], and EG3D [7] employ GAN-based 3D generators conditioned on latent vectors to produce category-specific textured 3D assets. Recently, as seen in GigaGAN [35], the Generative Adversarial Networks (GANs) methods are better suited for high-frequency details than diffusion models. Furthermore, IT3D [10] proposes a novel Diffusion-GAN dual training strategy to overcome the view inconsistency challenges. However, the training process of GAN is prone to the issue of mode collapse, which limits the diversity of generation results.
Vision Transformers. Computer vision has undergone a transformative shift with the emergence of vision transformers [16, 59]. Their ability to model global and local contexts has sparked interest in exploring methods to pre-train these backbones. MoCoV3 [9] and MAE [23] delved into different aspects of self-supervised Vision Transformers. In particular, MAE achieved state-of-the-art results by drawing inspiration from BERT [15], which involves randomly masking words in sentences, and using masked image reconstruction. Subsequent research extended this concept to the 3D domain [24, 39, 52, 77], leading to impressive results in downstream 3D vision tasks.
Figure 2: Scene123’s pipeline includes two key modules: the consistency-enhanced MAE and the 3D-aware generative refinement module. The former generates adjacent views from an input image via warping, using the MAE model to inpaint unseen areas with global semantics and optimizing an implicit neural field for viewpoint consistency. The latter generates realistic scene videos from the input image with a pre-trained video generation model, enhancing realism through adversarial loss with rendered images.
3 Methodology
3.1 Overview
In this paper, we propose a novel 3D scene generation framework based on one prompt, a single image or textual description, ensuring viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, as shown in Fig. 2. We first design the Consistency-Enchanced MAE module to fill unseen areas in novel views by injecting global semantic information in Sec. 3.2. With support views in the initialized database $\mathbf{S}$ generated by the Consistency-Enhanced MAE module, we employ a NeRF network to represent the 3D scene as the physical constraints of view consistency in Sec. 3.3. Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images in  3.4. Finally, the optimization and implementation details of the model are detailed in Sec. 3.5.
3.2 Consistency-Enhanced MAE Scene Completion
Scene Initialization. Given the reference image $\mathbf{I_{0}}$, notably, for only text prompt input, we utilize the stable diffusion model [63] to generate initial image $\mathbf{I_{0}}$, we then feed this image into the off-the-shelf depth estimation model [48], and take the output as a geometric prior for the target scene, denoted as $\mathbf{D_{0}}$. Inspired by [95], we construct an original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$ via the depth image-based rendering (DIBR) method [18], where $N$ denotes the number of initial viewpoints. Specifically, for each pixel $x$ in $\mathbf{I_{0}}$ and its depth value $y$ in $\mathbf{D_{0}}$, we compute its corresponding pixel $x_{0\rightarrow m}$ and depth $y_{0\rightarrow m}$ on a surrounding view $m$, $\left[x_{0\rightarrow m},y_{0\rightarrow m}\right]^{T}=\mathbf{K}\mathbf{P}_{m}\mathbf{P}_{0}^{-1}\mathbf{K}^{-1}[x,y]^{T}$, where $\mathbf{K}$ and $\mathbf{P}_{m}$ indicate the intrinsic matrix and the camera pose in view $m$. This database provides additional views and depth information, which could prevent the model from overfitting to the initial view.
MAE Scene Completion. However, the original database $\mathcal{S}_{0}$ will inevitably have missing content since the information in the initial scene is derived from the single image $\mathbf{I_{0}}$ to construct the initialized database $\mathcal{S}$. Directly applying the original database to the 3D scene representation would inevitably suffer from limited 3D consistency of the generated scenes [95, 28]. Inspired by  [23, 30, 94], we design a Consistency-Enhanced MAE module to effectively ensure the consistency of unseen areas across views. Specifically, we first design a discrete codebook to distribute global information to every invisible area in the original database. Specifically, we use a powerful pre-trained model on the ImageNet dataset [66]. We represent the codebook as =$\mathcal{E}=\{e_{1},e_{2},...,e_{N}\}\in\mathbb{R}^{N\times n_{q}}$ , where $N$ stands for the total count of prototype vectors, $n_{q}$ denotes the dimensionality of individual vectors, and $e_{i}$ symbolizes each specific embedding vector. Specifically, given the input image $x\in\mathbb{R}^{H\times W\times 3}$, the VQ-VAE [79] employs an encoder $E$ to extract a continuous feature representation: $\hat{z}=E(x)\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the height and width of the feature map. This continuous feature map $\hat{z}$ is then subjected to a quantization process $Q$, aligning it with its nearest codebook entry $e_{k}$ to obtain its discrete representation $z_{q}$, $z_{q}=Q(\hat{z}):=\underset{e_{k}\in\mathcal{E}}{\operatorname{argmin}}\left\|\hat{z}_{ij}-e_{k}\right\|_{2}$, where $\hat{z}_{ij}\in\mathbb{R}^{c}$.
With the pre-trained codebook, we then utilize the MAE encoder to encode the input images from the original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$, forming the image-conditional $\mathbf{s_{c}}=\{s_{1},s_{2},...,s_{M}\}$, each embedding vector $s_{i}$ queries the valuable prior information from the given codebook via the cross-attention mechanism, $\mathrm{Q}\leftarrow f_{\mathrm{q}}(\mathbf{s_{c}}),\mathrm{K}\leftarrow f_{\mathrm{k}}(\mathcal{E}),\mathrm{V}\leftarrow f_{\mathrm{v}}(\mathcal{E}),\mathbf{s_{u}}\leftarrow\text{Cross-Attention}(\mathrm{Q},\mathrm{K},\mathrm{V})=\text{Softmax}(\frac{\mathrm{Q}\mathrm{K}^{\mathrm{T}}}{\sqrt{d_{k}}})$, where $f_{q}$, $f_{k}$, and $f_{v}$ are the query, key, and value linear projections, respectively. Consequently, the global semantic information contained in the codebook is maintained to distribute global information to every invisible area. Then, we utilize the MAE decoder to decode the feature $\mathbf{s_{u}}$, deriving the initialized database $\mathcal{S}$, as shown in Fig. 2.
3.3 3D Scene Representation
With these support views in the initialized database $\mathcal{S}$, along with the initial view $\mathbf{I_{0}}$, we aim to generate 3D scenes through robust 3D scene representations to provide physical-level surface consistency constraints. Recently, 3D Gaussian Splatting (3DGS) [36] has recently emerged as an explicit alternative to NeRF [49], offering extremely fast rendering speeds and a memory-efficient backward pass. However, 3DGS occasionally have popping artifacts when our optimization creates large Gaussians, tending to happen in regions with view-dependent appearances. Moreover, Unlike explicit representations like polygon meshes or point clouds [61, 50], which are hard to represent complex geometry, NeRF shows its power in representing arbitrarily complex scenes. Therefore, we employ a NeRF network $f_{\theta}$ to represent the 3D scene. Specifically, in NeRF representation, volume rendering [49] is used to accumulate the color in the radiance fields, $\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\mathrm{d}t$, where $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ represents the 3D coordinates of sampled points on the camera ray emitted from the camera center $\mathbf{o}$ with the direction $\mathbf{d}$. $t_{n}$ and $t_{f}$ indicate the near and far sampling bounds. $(\mathbf{c},\sigma)=f_{\theta}(\mathbf{r}(t))$ are the predicted color and density of the sampled point along the ray. $T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\mathrm{d}s\right)$ is the accumulated transmittance. Different from NeRF that takes both the 3D coordinate $\mathbf{r}(t)$ and view direction $\mathbf{d}$ to predict the radiance $\mathbf{c}(\mathbf{r}(t),\mathbf{d})$, we omit $\mathbf{d}$ to avoid the effect of view-dependent specularity.
However, due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions [43]. Despite the design of the consistency-enhanced MAE module greatly improves the inconsistent between different views, the estimated depth rendered from NeRF still may be inconsistent between views. Inspired by  [95], we globally align these two depth maps by compensating for mean scale and value differences.
3.4 Video-assisted 3D-Aware Generative Refinement
Support Set Generation. Given the Reference image $\mathbf{I_{0}}$, we employ a image-to-video pipeline to generate a support set of enhanced quality, termed $D$, which is conditioned on the input reference image. For the image-to-video pipeline, we opt for Stable Video Diffusion(SVD) [4], which is trained to generate smooth and consistent videos on large-scale datasets of real and high-quality videos. The exposure to superior data quantity and quality makes it more generalizable and multi-view consistent, and the flexibility of the SVD architecture makes it amenable to be finetuned for camera controllability. In the context of the SVD image-to-video [4] pipeline, a noise-augmented [27] version of the conditioning frame channel-wise is concatenated to the input of the UNet [64]. In addition, the temporal attention layers in SVD naturally assist in the consistent multi-view generation without needing any explicit 3D structures like in [42].
3D-Aware Generative Refinement. The capabilities of Generative Adversarial Networks (GANs) shine in scenarios involving datasets characterized by high variance [7]. GANs have the ability to learn both geometry and texture-related knowledge from datasets, subsequently guiding the model to converge towards the same high-quality distribution exhibited by the generated support set. In our approach, we designate the video diffusion model as a generator. As shown in Fig. 3, given a reference image, SVD can generate consistent multi-view images at one time, constructing the generated support set. We then incorporate a discriminator initialized with random values. In this setup, the Support Set $D$ is treated as real data, while the renderings of the 3D neural radiance field model represent fake data. The role of the discriminator involves learning the distribution discrepancy between the renderings and $D$, subsequently contributing to the discrimination loss, which in turn updates the 3D neural radiance field model. This 3D-aware generative refinement module utilizes the discrimination loss $L_{\mathbf{dist}}$ to guide the updating direction and enhances the model’s ability to produce intricate geometry and texture details.
Figure 3: Data samples generated through image-to-video pipeline.
3.5 Optimization and Implementation Details
Optimization. In addition to discrimination loss, we also utilize a $L_{2}$ loss, depth loss and a transmittance loss to optimize the radiance field of the 3D scene, following previous NeRF-based works [8, 75, 95]. The RGB loss $L_{\mathbf{RGB}}$ is defined as a $L_{2}$ loss between the render pixel $\boldsymbol{C}^{R}$ and the color $\boldsymbol{C}$ generated by the MAE model. Different from previous works that employ regularized depth losses to handle uncertainty or scale-variant problems [62, 67], we adopt a stricter depth loss $L_{\mathbf{depth}}$ to minimize the $L_{2}$ distance between the rendered depth and the estimated depth. Moreover, we compute a depth-aware transmittance loss $L_{\mathbf{T}}$ [33, 95] to encourage the NeRF network to produce empty density before the camera ray reaches the expected depth $\mathbf{\hat{z}}$, $L_{\mathbf{T}}=\|\mathbf{T}(t)\cdot\mathbf{m}(t)\|_{2}$, where $\mathbf{m}(t)$ is a mask indicator that satisfies $\mathbf{m}(t)=1$ when $t<\mathbf{\hat{z}}$, otherwise $\mathbf{m}(t)=0$. $\mathbf{\hat{z}}$ is the pixel-wise depth value in the estimated depth map, and $\mathbf{T}(t)$ is the accumulated transmittance. Therefore, the total loss function is then defined as, $L_{\mathbf{total}}=L_{\mathbf{RGB}}+\lambda_{d}L_{\mathbf{depth}}+\lambda_{t}L_{\mathbf{T}}+\lambda_{dist}L_{\mathbf{dist}}$, where $\lambda_{d}$, $\lambda_{T}$, $\lambda_{dist}$ are constant hyperparameters balancing between depth loss, transmittance loss and discrimination loss.
Implementation Details. We implement the Scene123 with the Pytorch framework [53] and adopt TensoRF [8] as the radiance field. To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints. The dimension of the masked codebook is 2048$\times$16. For only text prompt input, we utilize the stable diffusion model in version 2.0 [63] to generated initial image $\mathbf{I_{0}}$ related to the input prompt. Moreover, for depth estimation, we use the boosting monocular depth estimation method [48] with pre-trained LeReS model [91] to estimate the depth for each view. For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD [4] but finetuned for 25 frame generation. During training, we use the same setting as [8] for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{d}=1000$, $\lambda_{dist}=0.0001$.
Figure 4: Qualitative results (zoom-in to view better) of methods capable of processing either a single image or text description. We both visualize the texture and depth from novel views within the scene.
4 Experiments
4.1 Experimental Setup
Dataset and baselines. Since the perpetual 3D scene generation is a new task without an existing dataset, we use real or generated high-quality images as input [93] for evaluation in our experiments. We consider two state-of-the-art 3D scene generation methods as our baselines, WonderJourney [93] and LucidDreamer [12] to compare the performance of the image or text prompt as a condition input for 3D scene generation. WonderJourney and LucidDreamer rely on an off-the-shelf general-purpose depth estimation model to project the hallucinated 2D scene extensions into a 3D representation. Specifically, WonderJourney designs a fully modularized model to generate sequences of 3D scenes. LucidDreamer utilizes Stable Diffusion [63] and 3D Gaussian splatting [36] to create diverse high-quality 3D scenes. For only text prompt input, we consider Text2NeRF [95], ProlificDreamer [86] and SJC [84] as the baselines, which perform well for the text to 3D generation. Text2NeRF generates NeRFs from text with the aid of text-to-image diffusion models to represent 3D scenes. SJC proposes a score distillation sampling (SDS) loss to extract deep semantic priors from pre-trained text-to-image diffusion models [63] and supervise the generative network of 3D models, and ProlificDreamer improves SDS by training an additional view-dependent Low-Rank Adaptation [29] for better 3D consistency.
Evaluation Metrics. A proficient 3D scene generation methodology should not only replicate the reference view closely but also preserve semantic and textural consistency when viewed from new perspectives. Therefore, to evaluate the overall quality of the generated 3D scenes, for the image prompt input, we choose the following metrics FID [25], CLIP similarity [58], and contextual distance [45]. S pecifically, we compute the FID score between the generated images and the input image, and lower scores have been shown to correlate well with higher quality images. Moreover, we measure CLIP similarity and contextual distance to evaluate the consistency of semantics and textures, respectively, between new perspectives and the reference view. For the text-driven 3D scene generation, we compute the CLIP similarity [58] between the text input and the generated views.
Figure 5: Qualitative results (zoom-in to view better) of methods that generate scenes only from textual input. We visualize the texture details from novel views within the scene.
4.2 Performance Comparison
As shown in Tab. 1, Tab. 2, we evaluate the quality of generated 3D scenes across baselines quantitatively. In Tab. 1, Our method achieves superior 3D consistency and high fidelity in generating 3D scenes, as evidenced by the best performance from FID, CLIP similarity, and contextual distance. Our method achieves higher CLIP similarity than text-driven 3D scene generation baselines in Tab. 2, showing the semantic relevance between the generated scene and the input text.
The qualitative results are drawn in Fig. 4 and Fig. 5. Our method can ensure more viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, giving a single image or textual description. Obviously, WonderJourney and LucidDreamer seem to generate 3D scenes with artifacts due to the estimation error, which relies on an off-the-shelf general-purpose depth estimation model. SJC exhibits the ability to generate object-centric scenes with a dreamlike style and ProlificDreamer can produce cloudy results. Moreover, Text2NeRF can suffer from noticeable artifacts and unrealism during the inpainting stage. Consequently, our method demonstrates superior qualitative performance compared to these baseline approaches.
Prompt WonderJourney[93] LucidDreamer[12] Ours
CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓
Image 88.201 1.539 156.619 91.034 1.918 156.651 93.565 1.538 89.829
Table 1: Quantitative comparison results of our method with the baseline WonderJourney and LucidDreamer, all capable of processing a single images or text. ↑ means the higher, the better, ↓ means the lower, the better.
Prompt WonderJourney[93] LucidDreamer[12] Text2NeRF[95] ProlificDreamer[86] SJC[84] Ours
CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑
Text 27.480 26.323 28.695 25.245 24.152 30.544
Table 2: Quantitative comparison results of our method with the baseline WonderJourney, LucidDreamer, Text2NeRF, ProlificDreamer, and SJC, all of which only accept text inputs. ↑ means the higher, the better, ↓ means the lower, the better.
4.3 Ablation Studies and Analysis
Effectiveness of the Consistency-Enhanced MAE. To verify the effectiveness of the Consistency-Enhanced MAE, we conduct ablation studies on different strategies: removing the MAE scene completion, denoted as w/o MAE; removing the masked VQ-VQE codebook, denoted w/o codebook; removing both MAE scene completion and the masked VQ-VAE codebook, denoted as w/o MAE&\&&codebook. As shown in Tab. 6 and Fig. 6, the integration of the consistency-enhanced MAE and the codebook significantly contributes to the generation of coherent 3D scenes. Fig. 6 demonstrates visually that when the model removes either of the MAE and codebook modules, it causes inconsistencies between the different views.
Method Image Text
CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
w/o MAE 88.284 1.722 124.242 28.598
w/o codebook 87.232 1.827 134.568 28.023
w/o MAE&\&&codebook 82.392 2.024 155.322 26.767
full model 93.565 1.538 89.829 30.544
[…]
tableAblation experiments regarding the MAE module and the codebook.
Figure 6: Results of the effectiveness of the MAE module and the codebook.
Effectiveness of the 3D-aware generative refinement module. To verify the effectiveness of the video-assisted 3D-aware generative refinement module, we conduct ablation studies on different strategies, including removing the discrimination loss, denoted as w/o GAN loss; replacing the real support set generated by the image-to-video pipeline with a set containing the same number of images using reference image duplication, denoted as w/o video-assisted. As shown in Tab. 7 and Fig. 7, the incorporation of the GAN-based training strategy significantly enhances the model’s ability to render detailed textures and complex geometries. The full model achieves the lowest FID values, while similar in other metrics, indicating the 3D-aware generative refinement module plays an important role in providing intricate geometry and texture details.
Method Image Text
CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
w/o GAN loss 92.324 1.563 156.244 30.422
w/o video-assisted 93.533 1.548 94.294 30.242
full model 93.565 1.538 89.829 30.544
[…]
tableAblation experiments regarding the 3D-aware generative refinement module.
Figure 7: Results of the effectiveness of the 3D-aware generative refinement module.
5 Conclusions and limitations
Conclusions. In this paper, we introduce Scene123, which surpasses existing 3D scene generation methods in scene consistency and realism, providing finer geometry and high-fidelity textures. Our method mainly relies on the consistency-enhanced MAE and the 3D-aware generative refinement module. The former exploits the inherent scene consistency constraints of implicit neural fields and integrates them with the MAE model with global semantics to inpaint adjacent views, ensuring viewpoint consistency. The latter uses the video generation model to produce realistic videos, enhancing the detail and realism of rendered views. With the help of these two modules, our method can generate high-quality 3D scenes from a single input prompt, whether real, virtual, or object-centered settings.
Limitations. Our method is both innovative and effective. However, the optimization process remains an area for potential enhancement, as it is currently constrained by our use of implicit neural fields to enforce physical consistency. In our future research, we aim to explore faster, more easily optimized, and sustainable three-dimensional representations of scenes. These will serve as mechanisms to articulate consistency surface constraints, thus accelerating the generation process.
Appendix A Appendix
A.1 Additional Implementation Details
A.1.1 GAN-based Training Strategy Details
For the incorporated discriminator, we adopt a similar architecture, regularization function, and loss weight as the EG3D model [7], with some distinctions. The vanilla structure of a 3D GAN involves a 3D generator that incorporates a super-resolution module, accompanied by a discriminator that accepts both coarse and fine image inputs [7]. Notably, due to the contextual disparities between text-to-3D and 3D GAN applications, we opt to omit the super-resolution component from the 3D GAN architecture. This choice stems from the consistent need to extract mesh or voxel representations from the 3D model within the context of text-to-3D.
A.2 Additional Qualitative Results
We provide additional qualitative results in Fig. 8, including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
Figure 8: More style visualization examples generated by Scene123. We provide additional results including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
A.3 User Study
For completeness, we follow previous works [86, 93] and conduct a user study by comparing Scene123 with WonderJourney [93], LucidDreamer [12] under 40 image prompts, 20 image prompts for each baseline. For text prompt input, we compare Scene123 with Text2NeRF [95], ProlificDreamer [86], SJC [84] under 60 text prompts, 20 text prompts for each baseline. The volunteers are shown the generated results of our Scene123 and baselines and asked to choose the better one in terms of fidelity, details and vividness. We collect results from 39 volunteers, yielding 3120 pairwise comparisons. The results are shown in Tab. 3. Our method outperforms all of the baselines.
Method WonderJourney LucidDreamer Text2NeRF ProlicDreamer SJC
Prefer baseline 31.12 23.42 46.27 10.25 5.50
Prefer Scene123 (Ours) 68.88 76.58 53.73 89.75 94.50
Table 3: Results of user study. The percentage of user preference (↑) is reported in the table.
A.4 Additional analysis
As depicted in Fig. 9, Scene123 doesn’t strictly require a high-quality support set. The novel views generated by our method are generally superior to the views generated by the image-to-video pipeline. This illustrates the adaptability of our approach to cope with such instances of generation failure.
Figure 9: Comparisons between support set generated by image-to-video pipeline and the generated views of Our method.

### v2

@@
 Abstract
-As Artificial Intelligence Generated Content (AIGC) advances, a variety of methods have been developed to generate text, images, videos, and 3D objects from single or multimodal inputs, contributing efforts to emulate human-like cognitive content creation.
+As Artificial Intelligence Generated Content (AIGC) advances, a variety of methods have been developed to generate text, images, videos, and 3D shapes from single or multimodal inputs, contributing efforts to emulate human-like cognitive content creation.
 However, generating realistic large-scale scenes from a single input presents a challenge due to the complexities involved in ensuring consistency across extrapolated views generated by models.
-Benefiting from recent video generation models and implicit neural representations, we propose Scene123, a 3D scene generation model, that not only ensures realism and diversity through the video generation framework but also uses implicit neural fields combined with Masked Autoencoders (MAE) to effectively ensures the consistency of unseen areas across views.
-Specifically, we initially warp the input image (or an image generated from text) to simulate adjacent views, filling the invisible areas with the MAE model.
-However, these filled images usually fail to maintain view consistency, thus we utilize the produced views to optimize a neural radiance field, enhancing geometric consistency.
+Benefiting from recent video generation models and implicit neural representations, we propose Scene123, a 3D scene generation model, which combines a video generation framework to ensure realism and diversity with implicit neural fields integrated with Masked Autoencoders (MAE) to effectively ensure the consistency of unseen areas across views.
+Specifically, the input image (or a text-generated image) is first warped to simulate adjacent views, with the invisible regions filled using the consistency-enhanced MAE model.
+Nonetheless, the synthesized images often exhibit inconsistencies in viewpoint alignment, thus we utilize the produced views to optimize a neural radiance field, enhancing geometric consistency.
 Moreover, to further enhance the details and texture fidelity of generated views, we employ a GAN-based Loss against images derived from the input image through the video generation model.
@@
 We show encourage video examples at https://yiyingyang12.github.io/Scene123.github.io/.
-Figure 1: Some examples generated by Scene123.
-For a single input image or text, our method can generate 3D scenes with consistent perspectives, fine geometry, and realistic textures, applicable to real, virtual, or object-centered scenes.
+Figure 1: Some examples generated by our Scene123.
+For a single input image or text, our method can generate 3D scenes with consistent views, fine geometry, and realistic textures, applicable to real, virtual, or object-centered scenes.
 1 Introduction
@@
 This is a fundamental issue in computer vision and graphics and a critical challenge in generative artificial intelligence.
-Recent advancements have demonstrated substantial progress through the use of vision-language models [58, 37], generative models such as Generative Adversarial Networks (GANs) [96], Variational autoencoders (VAEs) [67], or diffusion models [88], and scene representations like Neural Radiance Fields (NeRF) [49] or 3D Gaussian Splatting (3DGS) [36].
-Typically, these methods begin by generating images with a pretrained generative model [63, 90], or by directly using images for image-to-scene generation, then estimate additional 3D surface geometric details such as depth and normal [54, 38], and subsequently render the surface textures of the scenes using generative strategies like inpainting [85, 1].
-Moreover, some approaches update the geometric surfaces as new views are generated to maintain the coherence of the scene [95].
-However, these methods often rely on pre-trained models [58, 37], resulting in inconsistencies and artifacts in the generated scenes.
+Recent advancements have demonstrated substantial progress through the use of vision-language models (Radford et al.
+2021), generative models such as Generative Adversarial Networks (GANs) (Zhang et al.
+2019), Variational autoencoders (VAEs) (Sargent et al.
+2023a; Yang et al.
+2023b), or diffusion models (Yang et al.
+2023a), and scene representations like Neural Radiance Fields (NeRF) (Mildenhall et al.
+2021; Yin et al.
+2022; Ding et al.
+2023; Lu et al.
+2023; Yang et al.
+2024) or 3D Gaussian Splatting (3DGS) (Kerbl et al.
+2023).
+Typically, these methods begin by generating images with a pretrained generative model (Rombach et al.
+2022), or by directly using images for image-to-scene generation, then estimate additional 3D surface geometric details such as depth and normal (Piccinelli, Sakaridis, and Yu 2023; Li et al.
+2023; Yin et al.
+2023; Yin and Zhou 2020), and subsequently render the surface textures of the scenes using generative strategies like inpainting (Wang et al.
+2023b).
+Moreover, some approaches update the geometric surfaces as new views are generated to maintain the coherence of the scene (Zhang et al.
+2024b).
+However, these methods often rely on pre-trained models (Radford et al.
+2021; Li et al.
+2022), resulting in inconsistencies and artifacts in the generated scenes.
 Additionally, these methods face challenges in producing high-quality, coherent 3D representations across diverse and complex environments.
@@
 Conversely, flexibility requires the model to produce high-quality outputs under varying conditions, which may introduce inconsistencies in the generated scenes.
-It is noteworthy that scene synthesis based on multi-view images is a long-studied topic from the early Multi-View Stereo (MVS)  [70, 69] to recent implicit neural representations [92, 2], ensuring view consistency in generated scenes through multi-view matching mechanisms.
+It is noteworthy that scene synthesis based on multi-view images is a long-studied topic from the early Multi-View Stereo (MVS)  (Schönberger et al.
+2016; Schönberger and Frahm 2016) to recent implicit neural representations (Yu et al.
+2021), ensuring view consistency in generated scenes through multi-view matching mechanisms.
 However, when the input is reduced to a single image or a textual description, lacking reference multi-view images, the performance of these methods is greatly compromised.
-Fortunately, methodologies such as Masked Autoencoders (MAE)  [23] provide avenues for extrapolating to areas unseen in new views, while the incorporation of additional semantic layers augments the coherence of the synthesized scenes.
-Additionally, video generation models  [26, 44] facilitate the enhancement of scenes with richer priors and more detailed texture information.
+Fortunately, methodologies such as Masked Autoencoders (MAE)  (He et al.
+2022) provide avenues for extrapolating to areas unseen in new views, while the incorporation of additional semantic layers augments the coherence of the synthesized scenes.
+Additionally, video generation models  (Luo et al.
+2023) facilitate the enhancement of scenes with richer priors and more detailed texture information.
 Thus, how to utilize multi-view reconstruction with robust physical constraints alongside effective expansions in new viewpoints and video generation models is still an unreached area.
@@
 Subsequently, a consistency-enhanced MAE is designed to inpaint the unseen areas, with a shared codebook maintained to distribute global information to every invisible area.
-We implement a progressive strategy derived from Text2NeRF  [95] to incrementally update perspectives throughout this process.
+We implement a progressive strategy derived from Text2NeRF  (Zhang et al.
+2024b) to incrementally update perspectives throughout this process.
 Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images.
@@
 3) We introduce the video-assisted 3D-aware generative refinement module, which enhances scene reconstruction by integrating the diversity and realism of video generation models through a GAN-based function to significantly improve detail and texture fidelity.
-4) Extensive experiments validate the efficacy of Scene123, demonstrating greater accuracy in surface reconstruction, higher realism in reconstructed views, and better texture fidelity compared to the state-of-the-art methods.
-The code and weights will be available.
+4) Extensive experiments validate the efficacy of Scene123, demonstrating greater accuracy in surface reconstruction, higher realism in reconstructed views, and better texture fidelity compared to the SOTA methods.
+The data and code will be available.
 2 Related Work
-3D Scene Generation.
-Recent advancements in text-to-3D and image-to-3D scene generation have been remarkable.
-Most research in this domain can be classified into two main categories: object-centric [14, 34, 40, 41, 47, 56, 57, 60, 72, 74, 78, 86, 92], which focuses on objects without backgrounds, and holistic, which involves generating entire 3D scenes or 3D trajectories including backgrounds.
-Expanding these techniques to create fully unconstrained 3D scenes remains a significant challenge.
-Many recent advancements in text-driven 3D scene generation have focused on modeling 3D scenes using text inputs [32, 13, 95, 19, 28, 94, 73, 97].
-Due to the scarcity of paired text-3D scene data, most studies utilize Contrastive Language-Image Pre-training (CLIP) [58] or pre-trained text-to-image models to interpret the text input.
-Text2Scene [32] employs CLIP to model and stylize 3D scenes from text (or image) inputs by decomposing the scene into manageable sub-parts.
-Set-the-Scene [13] and Text2NeRF [95] generate NeRFs from text using text-to-image diffusion models to represent 3D scenes.
-SceneScape [19] and Text2Room [28] leverage a pre-trained monocular depth prediction model for enhanced geometric consistency and directly generate the 3D textured mesh representation of the scene.
+Text to 3D Scene Generation.
+Many recent advancements in text-driven 3D scene generation have focused on modeling 3D scenes using text inputs (Hwang, Kim, and Kim 2023; Zhang et al.
+2024b).
+Due to the scarcity of paired text-3D scene data, most studies utilize Contrastive Language-Image Pre-training (CLIP) (Radford et al.
+2021) or pre-trained text-to-image models to interpret the text input.
+Text2Scene (Hwang, Kim, and Kim 2023) employs CLIP to model and stylize 3D scenes from text (or image) inputs by decomposing the scene into manageable sub-parts.
+Set-the-Scene (Cohen-Bar et al.
+2023) and Text2NeRF (Zhang et al.
+2024b) generate NeRFs from text using text-to-image diffusion models to represent 3D scenes.
+SceneScape (Fridman et al.
+2024) and Text2Room (Höllein et al.
+2023) leverage a pre-trained monocular depth prediction model for enhanced geometric consistency and directly generate the 3D textured mesh representation of the scene.
 However, these methods depend on inpainted images for scene completion, which, while producing realistic visuals, suffer from limited 3D consistency.
-More recent studies [20, 2, 80] have successfully generated multi-object compositional 3D scenes.
-Another class of methods utilizes auxiliary inputs, such as layouts [55, 17], to enhance scene generation.
+More recent studies (Bai et al.
+2023) have successfully generated multi-object compositional 3D scenes.
+Another class of methods utilizes auxiliary inputs, such as layouts (Po and Wetzstein 2023), to enhance scene generation.
 Unlike text captions, which can be rather vague, some approaches generate 3D scenes from image inputs, where the 3D scene corresponds closely to the depicted image.
-Early scene generation methods often require specific scene data for training to obtain category-specific scene generators, such as GAUDI [3], or focus on single scene reconstruction based on the input image, such as PixelSynth [61] and Worldsheet [31].
+Early scene generation methods often require specific scene data for training to obtain category-specific scene generators, such as GAUDI (Bautista et al.
+2022), or focus on single scene reconstruction based on the input image, such as PixelSynth (Rockwell, Fouhey, and Johnson 2021) and Worldsheet (Hu et al.
+2021).
 However, these methods are often limited by the quality of the generation or the extensibility of the scene.
+Image to 3D Scene Generation.
 Recently, numerous studies have concentrated on generating 3D scenes from image inputs.
-PERF [83] generates 3D scenes from a single panoramic image, using diffusion models to supplement shadow areas.
-ZeroNVS [68] extends this capability by reconstructing both objects and environments in 3D from a single image.
+PERF (Wang et al.
+2023a) generates 3D scenes from a single panoramic image, using diffusion models to supplement shadow areas.
+ZeroNVS (Sargent et al.
+2023b) extends this capability by reconstructing both objects and environments in 3D from a single image.
 Despite its environmental reconstruction lacking some detail, the algorithm demonstrates an understanding of environmental contexts.
-LucidDreamer [12] and WonderJourney [93] utilize general-purpose depth estimation models to project hallucinated 2D scene extensions into 3D representations.
+LucidDreamer (Chung et al.
+2023) and WonderJourney (Yu et al.
+2023) utilize general-purpose depth estimation models to project hallucinated 2D scene extensions into 3D representations.
 However, these methods still face challenges in achieving realism, often producing artifacts due to the reliance on pre-trained models.
 Video Diffusion Models and 3D-aware GANs.
-Diffusion models [76] have recently emerged as powerful generative models capable of producing a wide array of images [5, 63, 65] and videos [4, 22, 81] by iteratively denoising a noise sample.
-Among these models, the publicly available Stable Diffusion (SD) [63] and Stable Video Diffusion (SVD) [4] exhibit strong generalization capabilities due to training on extremely large datasets such as LAION [71] and LVD [4].
+Diffusion models (Song et al.
+2020) have recently emerged as powerful generative models capable of producing a wide array of images (Blattmann et al.
+2023b) and videos (Blattmann et al.
+2023a) by iteratively denoising a noise sample.
+Among these models, the publicly available Stable Diffusion (SD) (Rombach et al.
+2022) and Stable Video Diffusion (SVD) (Blattmann et al.
+2023a) exhibit strong generalization capabilities due to training on extremely large datasets such as LAION (Schuhmann et al.
+2022) and LVD (Blattmann et al.
+2023a).
 Consequently, they are frequently used as foundational models for various generation tasks, including novel view synthesis.
 To enhance generalization and multi-view consistency, some contemporary works leverage the temporal priors in video diffusion models for object-centric 3D generation.
-For instance, IM-3D [46] and SV3D [82] explore the capabilities of video diffusion models in object-centric multi-view generation.
-V3D [11] extends this approach to scene-level novel view synthesis.
+For instance, IM-3D (Melas-Kyriazi et al.
+2024) and SV3D (Voleti et al.
+2024) explore the capabilities of video diffusion models in object-centric multi-view generation.
+V3D (Chen et al.
+2024b) extends this approach to scene-level novel view synthesis.
 However, these methods often produce unsatisfactory results for complex objects or scenes, leading to inconsistencies among multiple views or unrealistic geometries.
-Early works, like 3D-GAN [87], Pointflow [89], and ShapeRF [6] focus more on the category-specific texture-less geometric shape generation based on the representations of voxels or point clouds.
+Early works, like 3D-GAN (Wu et al.
+2016), Pointflow (Yang et al.
+2019), and ShapeRF (Cai et al.
+2020) focus more on the category-specific texture-less geometric shape generation based on the representations of voxels or point clouds.
 However, limited by the generation capabilities of GANs, these methods can only generate rough 3D assets of specific categories.
-Subsequently, HoloGAN [51], GET3D [21], and EG3D [7] employ GAN-based 3D generators conditioned on latent vectors to produce category-specific textured 3D assets.
-Recently, as seen in GigaGAN [35], the Generative Adversarial Networks (GANs) methods are better suited for high-frequency details than diffusion models.
-Furthermore, IT3D [10] proposes a novel Diffusion-GAN dual training strategy to overcome the view inconsistency challenges.
+Subsequently, HoloGAN (Nguyen-Phuoc et al.
+2019), GET3D (Gao et al.
+2022), and EG3D (Chan et al.
+2022) employ GAN-based 3D generators conditioned on latent vectors to produce category-specific textured 3D assets.
+Recently, as seen in GigaGAN (Kang et al.
+2023), the Generative Adversarial Networks (GANs) methods are better suited for high-frequency details than diffusion models.
+Furthermore, IT3D (Chen et al.
+2024a) proposes a novel Diffusion-GAN dual training strategy to overcome the view inconsistency challenges.
 However, the training process of GAN is prone to the issue of mode collapse, which limits the diversity of generation results.
-Vision Transformers.
-Computer vision has undergone a transformative shift with the emergence of vision transformers [16, 59].
-Their ability to model global and local contexts has sparked interest in exploring methods to pre-train these backbones.
-MoCoV3 [9] and MAE [23] delved into different aspects of self-supervised Vision Transformers.
-In particular, MAE achieved state-of-the-art results by drawing inspiration from BERT [15], which involves randomly masking words in sentences, and using masked image reconstruction.
-Subsequent research extended this concept to the 3D domain [24, 39, 52, 77], leading to impressive results in downstream 3D vision tasks.
 Figure 2: Scene123’s pipeline includes two key modules: the consistency-enhanced MAE and the 3D-aware generative refinement module.
 The former generates adjacent views from an input image via warping, using the MAE model to inpaint unseen areas with global semantics and optimizing an implicit neural field for viewpoint consistency.
-The latter generates realistic scene videos from the input image with a pre-trained video generation model, enhancing realism through adversarial loss with rendered images.
+The latter generates realistic videos from the input image with a pre-trained video generation model, enhancing realism through adversarial loss with rendered images.
 3 Methodology
@@
 2.
-We first design the Consistency-Enchanced MAE module to fill unseen areas in novel views by injecting global semantic information in Sec.
-3.2.
-With support views in the initialized database $\mathbf{S}$ generated by the Consistency-Enhanced MAE module, we employ a NeRF network to represent the 3D scene as the physical constraints of view consistency in Sec.
-3.3.
-Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images in  3.4.
-Finally, the optimization and implementation details of the model are detailed in Sec.
-3.5.
+We first design the Consistency-Enchanced MAE module to fill unseen areas in novel views by injecting global semantic information.
+With support views in the initialized database $\mathbf{S}$ generated by the Consistency-Enhanced MAE module, we employ a NeRF network to represent the 3D scene as the physical constraints of view consistency.
+Furthermore, to enhance the scene’s detail and realism, we employ the latest video generation technology, generating high-quality scene videos based on input images and performing adversarial enhancements with the rendered images.
+Finally, the optimization and implementation details of the model are detailed.
 3.2 Consistency-Enhanced MAE Scene Completion
 Scene Initialization.
-Given the reference image $\mathbf{I_{0}}$, notably, for only text prompt input, we utilize the stable diffusion model [63] to generate initial image $\mathbf{I_{0}}$, we then feed this image into the off-the-shelf depth estimation model [48], and take the output as a geometric prior for the target scene, denoted as $\mathbf{D_{0}}$.
-Inspired by [95], we construct an original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$ via the depth image-based rendering (DIBR) method [18], where $N$ denotes the number of initial viewpoints.
+Given the reference image $\mathbf{I_{0}}$, notably, for only text prompt input, we utilize the stable diffusion model (Rombach et al.
+2022) to generate initial image $\mathbf{I_{0}}$, we then feed this image into the off-the-shelf depth estimation model (Miangoleh et al.
+2021), and take the output as a geometric prior for the target scene, denoted as $\mathbf{D_{0}}$.
+Inspired by (Zhang et al.
+2024b), we construct an original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$ via the depth image-based rendering (DIBR) method (Fehn 2004), where $N$ denotes the number of initial viewpoints.
 Specifically, for each pixel $x$ in $\mathbf{I_{0}}$ and its depth value $y$ in $\mathbf{D_{0}}$, we compute its corresponding pixel $x_{0\rightarrow m}$ and depth $y_{0\rightarrow m}$ on a surrounding view $m$, $\left[x_{0\rightarrow m},y_{0\rightarrow m}\right]^{T}=\mathbf{K}\mathbf{P}_{m}\mathbf{P}_{0}^{-1}\mathbf{K}^{-1}[x,y]^{T}$, where $\mathbf{K}$ and $\mathbf{P}_{m}$ indicate the intrinsic matrix and the camera pose in view $m$.
@@
 However, the original database $\mathcal{S}_{0}$ will inevitably have missing content since the information in the initial scene is derived from the single image $\mathbf{I_{0}}$ to construct the initialized database $\mathcal{S}$.
-Directly applying the original database to the 3D scene representation would inevitably suffer from limited 3D consistency of the generated scenes [95, 28].
-Inspired by  [23, 30, 94], we design a Consistency-Enhanced MAE module to effectively ensure the consistency of unseen areas across views.
+Directly applying the original database to the 3D scene representation would inevitably suffer from limited 3D consistency of the generated scenes (Zhang et al.
+2024b; Höllein et al.
+2023).
+Inspired by  (He et al.
+2022; Hu et al.
+2023; Zhang et al.
+2024a), we design a Consistency-Enhanced MAE module to effectively ensure the consistency of unseen areas across views.
 Specifically, we first design a discrete codebook to distribute global information to every invisible area in the original database.
-Specifically, we use a powerful pre-trained model on the ImageNet dataset [66].
 We represent the codebook as =$\mathcal{E}=\{e_{1},e_{2},...,e_{N}\}\in\mathbb{R}^{N\times n_{q}}$ , where $N$ stands for the total count of prototype vectors, $n_{q}$ denotes the dimensionality of individual vectors, and $e_{i}$ symbolizes each specific embedding vector.
-Specifically, given the input image $x\in\mathbb{R}^{H\times W\times 3}$, the VQ-VAE [79] employs an encoder $E$ to extract a continuous feature representation: $\hat{z}=E(x)\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the height and width of the feature map.
-This continuous feature map $\hat{z}$ is then subjected to a quantization process $Q$, aligning it with its nearest codebook entry $e_{k}$ to obtain its discrete representation $z_{q}$, $z_{q}=Q(\hat{z}):=\underset{e_{k}\in\mathcal{E}}{\operatorname{argmin}}\left\|\hat{z}_{ij}-e_{k}\right\|_{2}$, where $\hat{z}_{ij}\in\mathbb{R}^{c}$.
-With the pre-trained codebook, we then utilize the MAE encoder to encode the input images from the original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$, forming the image-conditional $\mathbf{s_{c}}=\{s_{1},s_{2},...,s_{M}\}$, each embedding vector $s_{i}$ queries the valuable prior information from the given codebook via the cross-attention mechanism, $\mathrm{Q}\leftarrow f_{\mathrm{q}}(\mathbf{s_{c}}),\mathrm{K}\leftarrow f_{\mathrm{k}}(\mathcal{E}),\mathrm{V}\leftarrow f_{\mathrm{v}}(\mathcal{E}),\mathbf{s_{u}}\leftarrow\text{Cross-Attention}(\mathrm{Q},\mathrm{K},\mathrm{V})=\text{Softmax}(\frac{\mathrm{Q}\mathrm{K}^{\mathrm{T}}}{\sqrt{d_{k}}})$, where $f_{q}$, $f_{k}$, and $f_{v}$ are the query, key, and value linear projections, respectively.
+Specifically, given the input image $x\in\mathbb{R}^{H\times W\times 3}$, the VQ-VAE (Van Den Oord, Vinyals et al.
+2017) employs an encoder $E$ to extract a continuous feature representation: $\hat{z}=E(x)\in\mathbb{R}^{h\times w\times c}$, where $h$ and $w$ are the height and width of the feature map.
+This continuous feature map $\hat{z}$ is then subjected to a quantization process $Q$, aligning it with its nearest codebook entry $e_{k}$ to obtain its discrete representation $z_{q}$ as follows:
+$z_{q}=Q(\hat{z}):=\underset{e_{k}\in\mathcal{E}}{\operatorname{argmin}}\left\|\hat{z}_{ij}-e_{k}\right\|_{2},$ (1)
+where $\hat{z}_{ij}\in\mathbb{R}^{c}$.
+We pre-trained VQVAE on the ImageNet (Russakovsky et al.
+2015) dataset to equip the codebook with a more representative and generalized feature set.
+Subsequently, we fine-tuned it on specific scenes to better capture and represent the unique characteristics of each scene.
+Moreover, we then utilize the MAE encoder to encode the input images from the original database $\mathcal{S}_{0}=\left\{\left(\mathbf{I}_{i},\mathbf{D}_{i}\right)\right\}_{i=1}^{N}$, forming the image-conditional $\mathbf{s_{c}}=\{s_{1},s_{2},...,s_{M}\}$, each embedding vector $s_{i}$ queries the valuable prior information from the given codebook via the cross-attention mechanism,
+$\begin{split}\mathrm{Q}\leftarrow f_{\mathrm{q}}(\mathbf{s_{c}}),\qquad\enspace\mathrm{K}\leftarrow f_{\mathrm{k}}(\mathcal{E}),\qquad\enspace\mathrm{V}\leftarrow f_{\mathrm{v}}(\mathcal{E})\\ \mathbf{s_{u}}\leftarrow\text{Cross-Attention}(\mathrm{Q},\mathrm{K},\mathrm{V})=\text{Softmax}(\frac{\mathrm{Q}\mathrm{K}^{\mathrm{T}}}{\sqrt{d_{k}}}),\end{split}$ (2)
+where $f_{q}$, $f_{k}$, and $f_{v}$ are the query, key, and value linear projections, respectively.
 Consequently, the global semantic information contained in the codebook is maintained to distribute global information to every invisible area.
@@
 2.
+Using global semantics as Key and Value in cross-attention allows the model to integrate comprehensive scene information, maintaining coherence.
+This approach enhances multi-view consistency by providing a unified understanding of the scene, reducing artifacts, and ensuring seamless integration of novel views.
 3.3 3D Scene Representation
 With these support views in the initialized database $\mathcal{S}$, along with the initial view $\mathbf{I_{0}}$, we aim to generate 3D scenes through robust 3D scene representations to provide physical-level surface consistency constraints.
-Recently, 3D Gaussian Splatting (3DGS) [36] has recently emerged as an explicit alternative to NeRF [49], offering extremely fast rendering speeds and a memory-efficient backward pass.
-However, 3DGS occasionally have popping artifacts when our optimization creates large Gaussians, tending to happen in regions with view-dependent appearances.
-Moreover, Unlike explicit representations like polygon meshes or point clouds [61, 50], which are hard to represent complex geometry, NeRF shows its power in representing arbitrarily complex scenes.
-Therefore, we employ a NeRF network $f_{\theta}$ to represent the 3D scene.
-Specifically, in NeRF representation, volume rendering [49] is used to accumulate the color in the radiance fields, $\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\mathrm{d}t$, where $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ represents the 3D coordinates of sampled points on the camera ray emitted from the camera center $\mathbf{o}$ with the direction $\mathbf{d}$.
+In this work, we employ a NeRF network $f_{\theta}$ to represent the 3D scene.
+Specifically, in NeRF representation, volume rendering (Mildenhall et al.
+2021) is used to accumulate the color in the radiance fields,
+$\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)\sigma(\mathbf{r}(t))\mathbf{c}(\mathbf{r}(t),\mathbf{d})\mathrm{d}t,$ (3)
+where $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ represents the 3D coordinates of sampled points on the camera ray emitted from the camera center $\mathbf{o}$ with the direction $\mathbf{d}$.
 $t_{n}$ and $t_{f}$ indicate the near and far sampling bounds.
 $(\mathbf{c},\sigma)=f_{\theta}(\mathbf{r}(t))$ are the predicted color and density of the sampled point along the ray.
-$T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\mathrm{d}s\right)$ is the accumulated transmittance.
+$T(t)=\exp\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(s))\mathrm{d}s\right),$ (4)
+where $T(t)$ is the accumulated transmittance.
 Different from NeRF that takes both the 3D coordinate $\mathbf{r}(t)$ and view direction $\mathbf{d}$ to predict the radiance $\mathbf{c}(\mathbf{r}(t),\mathbf{d})$, we omit $\mathbf{d}$ to avoid the effect of view-dependent specularity.
-However, due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions [43].
-Despite the design of the consistency-enhanced MAE module greatly improves the inconsistent between different views, the estimated depth rendered from NeRF still may be inconsistent between views.
-Inspired by  [95], we globally align these two depth maps by compensating for mean scale and value differences.
+However, due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions (Luo et al.
+2020).
+Despite the design of the consistency-enhanced MAE module greatly improving the inconsistency between different views, the estimated depth rendered from NeRF still may be inconsistent between views.
+Inspired by Text2NeRF (Zhang et al.
+2024b), we globally align these two depth maps by compensating for mean scale and value differences.
+Specifically, we first we first perform global alignment by calculating the average $s$ and depth offset $\delta$ to approximate the mean scale and value differences, and then we finetune a pre-trained depth alignment network to produce a locally aligned depth map.
+More details will be shown in the Appendix.
+Figure 3: Qualitative results (zoom-in to view better) of methods capable of processing a single image prompt. We both visualize the texture and depth from novel views within the scene.
 3.4 Video-assisted 3D-Aware Generative Refinement
@@
 Given the Reference image $\mathbf{I_{0}}$, we employ a image-to-video pipeline to generate a support set of enhanced quality, termed $D$, which is conditioned on the input reference image.
-For the image-to-video pipeline, we opt for Stable Video Diffusion(SVD) [4], which is trained to generate smooth and consistent videos on large-scale datasets of real and high-quality videos.
+For the image-to-video pipeline, we opt for Stable Video Diffusion(SVD) (Blattmann et al.
+2023a), which is trained to generate smooth and consistent videos on large-scale datasets of real and high-quality videos.
 The exposure to superior data quantity and quality makes it more generalizable and multi-view consistent, and the flexibility of the SVD architecture makes it amenable to be finetuned for camera controllability.
-In the context of the SVD image-to-video [4] pipeline, a noise-augmented [27] version of the conditioning frame channel-wise is concatenated to the input of the UNet [64].
-In addition, the temporal attention layers in SVD naturally assist in the consistent multi-view generation without needing any explicit 3D structures like in [42].
+In the context of the SVD image-to-video (Blattmann et al.
+2023a) pipeline, a noise-augmented (Ho et al.
+2022) version of the conditioning frame channel-wise is concatenated to the input of the UNet (Ronneberger, Fischer, and Brox 2015).
+In addition, the temporal attention layers in SVD naturally assist in the consistent multi-view generation without needing any explicit 3D structures like in (Liu et al.
+2023).
 3D-Aware Generative Refinement.
-The capabilities of Generative Adversarial Networks (GANs) shine in scenarios involving datasets characterized by high variance [7].
+The capabilities of Generative Adversarial Networks (GANs) shine in scenarios involving datasets characterized by high variance (Chan et al.
+2022).
 GANs have the ability to learn both geometry and texture-related knowledge from datasets, subsequently guiding the model to converge towards the same high-quality distribution exhibited by the generated support set.
@@
 As shown in Fig.
-3, given a reference image, SVD can generate consistent multi-view images at one time, constructing the generated support set.
+4, given a reference image, SVD can generate consistent multi-view images at one time, constructing the generated support set.
 We then incorporate a discriminator initialized with random values.
@@
 The role of the discriminator involves learning the distribution discrepancy between the renderings and $D$, subsequently contributing to the discrimination loss, which in turn updates the 3D neural radiance field model.
-This 3D-aware generative refinement module utilizes the discrimination loss $L_{\mathbf{dist}}$ to guide the updating direction and enhances the model’s ability to produce intricate geometry and texture details.
-Figure 3: Data samples generated through image-to-video pipeline.
+This 3D-aware generative refinement module utilizes the discrimination loss $L_{\mathbf{dist}}$, which can help guide the updating direction and enhance the model’s ability to produce intricate geometry and texture details.
+Figure 4: Data samples generated via image-to-video model.
+The camera motion in the SVD model is sometimes limited(Blattmann et al.
+2023a), making it unsuitable for directly training the NeRF model.
+Instead, we use our 3D-Aware Generative Refinement approach, which leverages a discriminator to optimize the process.
+Despite the small range of camera motion in the images generated by SVD, the adversarial training process provides necessary regularization, demonstrating that even with limited camera motion, the discrimination loss remains effective and contributes to generating high-quality 3D scenes.
+Figure 5: Qualitative results (zoom-in to view better) of methods that generate scenes from textual input. We both visualize the texture and depth from novel views within the scene.
 3.5 Optimization and Implementation Details
-Optimization.
-In addition to discrimination loss, we also utilize a $L_{2}$ loss, depth loss and a transmittance loss to optimize the radiance field of the 3D scene, following previous NeRF-based works [8, 75, 95].
+In addition to discrimination loss, we also utilize a $L_{2}$ loss, depth loss, and a transmittance loss to optimize the radiance field of the 3D scene, following previous NeRF-based works (Chen et al.
+2022; Song et al.
+2023; Zhang et al.
+2024b).
 The RGB loss $L_{\mathbf{RGB}}$ is defined as a $L_{2}$ loss between the render pixel $\boldsymbol{C}^{R}$ and the color $\boldsymbol{C}$ generated by the MAE model.
-Different from previous works that employ regularized depth losses to handle uncertainty or scale-variant problems [62, 67], we adopt a stricter depth loss $L_{\mathbf{depth}}$ to minimize the $L_{2}$ distance between the rendered depth and the estimated depth.
-Moreover, we compute a depth-aware transmittance loss $L_{\mathbf{T}}$ [33, 95] to encourage the NeRF network to produce empty density before the camera ray reaches the expected depth $\mathbf{\hat{z}}$, $L_{\mathbf{T}}=\|\mathbf{T}(t)\cdot\mathbf{m}(t)\|_{2}$, where $\mathbf{m}(t)$ is a mask indicator that satisfies $\mathbf{m}(t)=1$ when $t<\mathbf{\hat{z}}$, otherwise $\mathbf{m}(t)=0$.
+Different from previous works that employ regularized depth losses to handle uncertainty or scale-variant problems (Roessle et al.
+2022; Sargent et al.
+2023a), we adopt a stricter depth loss $L_{\mathbf{depth}}$ to minimize the $L_{2}$ distance between the rendered depth and the estimated depth.
+Moreover, we compute a depth-aware transmittance loss $L_{\mathbf{T}}$ (Jain et al.
+2022; Zhang et al.
+2024b) to encourage the NeRF network to produce empty density before the camera ray reaches the expected depth $\mathbf{\hat{z}}$, $L_{\mathbf{T}}=\|\mathbf{T}(t)\cdot\mathbf{m}(t)\|_{2}$, where $\mathbf{m}(t)$ is a mask indicator that satisfies $\mathbf{m}(t)=1$ when $t<\mathbf{\hat{z}}$, otherwise $\mathbf{m}(t)=0$.
 $\mathbf{\hat{z}}$ is the pixel-wise depth value in the estimated depth map, and $\mathbf{T}(t)$ is the accumulated transmittance.
-Therefore, the total loss function is then defined as, $L_{\mathbf{total}}=L_{\mathbf{RGB}}+\lambda_{d}L_{\mathbf{depth}}+\lambda_{t}L_{\mathbf{T}}+\lambda_{dist}L_{\mathbf{dist}}$, where $\lambda_{d}$, $\lambda_{T}$, $\lambda_{dist}$ are constant hyperparameters balancing between depth loss, transmittance loss and discrimination loss.
+Therefore, the total loss function is then defined as,
+$L_{\mathbf{total}}=L_{\mathbf{RGB}}+\lambda_{d}L_{\mathbf{depth}}+\lambda_{t}L_{\mathbf{T}}+\lambda_{dist}L_{\mathbf{dist}},$ (5)
+where $\lambda_{d}$, $\lambda_{T}$, $\lambda_{dist}$ are constant hyperparameters balancing depth, transmittance, and discrimination losses.
 Implementation Details.
-We implement the Scene123 with the Pytorch framework [53] and adopt TensoRF [8] as the radiance field.
+We implement our Scene123 with the Pytorch framework (Paszke et al.
+2019) and adopt TensoRF (Chen et al.
+2022) as the radiance field.
 To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints.
 The dimension of the masked codebook is 2048$\times$16.
-For only text prompt input, we utilize the stable diffusion model in version 2.0 [63] to generated initial image $\mathbf{I_{0}}$ related to the input prompt.
-Moreover, for depth estimation, we use the boosting monocular depth estimation method [48] with pre-trained LeReS model [91] to estimate the depth for each view.
-For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD [4] but finetuned for 25 frame generation.
-During training, we use the same setting as [8] for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{d}=1000$, $\lambda_{dist}=0.0001$.
-Figure 4: Qualitative results (zoom-in to view better) of methods capable of processing either a single image or text description.
-We both visualize the texture and depth from novel views within the scene.
+For only text prompt input, we utilize the stable diffusion model in version 2.0 (Rombach et al.
+2022) to generated initial image $\mathbf{I_{0}}$ related to the input prompt.
+Moreover, for depth estimation, we use the boosting monocular depth estimation method (Miangoleh et al.
+2021) with pre-trained LeReS model (Yin et al.
+2021) to estimate the depth for each view.
+For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD (Blattmann et al.
+2023a) but finetuned for 25 frame generation.
+During training, we use the same setting as (Chen et al.
+2022) for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{t}=0.001$, $\lambda_{dist}=0.001$.
 4 Experiments
@@
 Dataset and baselines.
-Since the perpetual 3D scene generation is a new task without an existing dataset, we use real or generated high-quality images as input [93] for evaluation in our experiments.
-We consider two state-of-the-art 3D scene generation methods as our baselines, WonderJourney [93] and LucidDreamer [12] to compare the performance of the image or text prompt as a condition input for 3D scene generation.
+Since the perpetual 3D scene generation is a new task without an existing dataset, we use real or generated high-quality images as input (Yu et al.
+2023) for evaluation in our experiments.
+We consider two state-of-the-art 3D scene generation methods as our baselines, WonderJourney (Yu et al.
+2023) and LucidDreamer (Chung et al.
+2023) to compare the performance of the image or text prompt as a condition input for 3D scene generation.
 WonderJourney and LucidDreamer rely on an off-the-shelf general-purpose depth estimation model to project the hallucinated 2D scene extensions into a 3D representation.
 Specifically, WonderJourney designs a fully modularized model to generate sequences of 3D scenes.
-LucidDreamer utilizes Stable Diffusion [63] and 3D Gaussian splatting [36] to create diverse high-quality 3D scenes.
-For only text prompt input, we consider Text2NeRF [95], ProlificDreamer [86] and SJC [84] as the baselines, which perform well for the text to 3D generation.
+LucidDreamer utilizes Stable Diffusion (Rombach et al.
+2022) and 3D Gaussian splatting (Kerbl et al.
+2023) to create diverse high-quality 3D scenes.
+For text prompts, besides WonderJourney and LucidDreamer, we include Text2NeRF (Zhang et al.
+2024b) as the baseline, which performs well for the text-to-3D generation.
 Text2NeRF generates NeRFs from text with the aid of text-to-image diffusion models to represent 3D scenes.
-SJC proposes a score distillation sampling (SDS) loss to extract deep semantic priors from pre-trained text-to-image diffusion models [63] and supervise the generative network of 3D models, and ProlificDreamer improves SDS by training an additional view-dependent Low-Rank Adaptation [29] for better 3D consistency.
+Method Input Visual Quality Optimization Time (hours)
+CLIP-Similarity↑ BRISQUE↓ NIQE↓
+WonderJourney (Yu et al. 2023) Image&Text 27.480 67.012 12.022 0.208
+LucidDreamer (Chung et al. 2023) Image&Text 26.663 46.266 6.652 0.220
+Text2NeRF (Zhang et al. 2024b) Text 28.695 24.498 4.618 1.525
+Scene123 (Ours) Image&Text 30.544 20.324 2.522 1.433
+Table 1: Quantitative comparison results of our method with the baseline WonderJourney, LucidDreamer and Text2NeRF ↑ means the higher, the better, ↓ means the lower, the better.
 Evaluation Metrics.
-A proficient 3D scene generation methodology should not only replicate the reference view closely but also preserve semantic and textural consistency when viewed from new perspectives.
-Therefore, to evaluate the overall quality of the generated 3D scenes, for the image prompt input, we choose the following metrics FID [25], CLIP similarity [58], and contextual distance [45].
-S pecifically, we compute the FID score between the generated images and the input image, and lower scores have been shown to correlate well with higher quality images.
-Moreover, we measure CLIP similarity and contextual distance to evaluate the consistency of semantics and textures, respectively, between new perspectives and the reference view.
-For the text-driven 3D scene generation, we compute the CLIP similarity [58] between the text input and the generated views.
-Figure 5: Qualitative results (zoom-in to view better) of methods that generate scenes only from textual input. We visualize the texture details from novel views within the scene.
-4.2 Performance Comparison
+Following Text2NeRF, we evaluate the quality of our generated images using CLIP Score (CLIP-Similarity), Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) (Mittal, Moorthy, and Bovik 2012) and Natural Image Quality Evaluator (NQIE) (Mittal, Soundararajan, and Bovik 2012).
+4.2 Performance Comparisons
 As shown in Tab.
-1, Tab.
-2, we evaluate the quality of generated 3D scenes across baselines quantitatively.
-In Tab.
-1, Our method achieves superior 3D consistency and high fidelity in generating 3D scenes, as evidenced by the best performance from FID, CLIP similarity, and contextual distance.
-Our method achieves higher CLIP similarity than text-driven 3D scene generation baselines in Tab.
-2, showing the semantic relevance between the generated scene and the input text.
+1, we evaluate the quality of generated 3D scenes across baselines quantitatively and report the average evaluation scores of CLIP-Similarity BRISQUE and NIQE for the generated images produced by different methods.
+Clearly, our method surpasses the baselines by generating higher-quality 3D scenes, as indicated by lower BRISQUE and NIQE values.
+Moreover, our method ensures the semantic relevance between the generated scene and the input text, resulting in a higher CLIP score.
 The qualitative results are drawn in Fig.
@@
 Our method can ensure more viewpoint consistency and realistic surface textures for both real and synthetically styled scenes, giving a single image or textual description.
-Obviously, WonderJourney and LucidDreamer seem to generate 3D scenes with artifacts due to the estimation error, which relies on an off-the-shelf general-purpose depth estimation model.
-SJC exhibits the ability to generate object-centric scenes with a dreamlike style and ProlificDreamer can produce cloudy results.
-Moreover, Text2NeRF can suffer from noticeable artifacts and unrealism during the inpainting stage.
+Obviously, Fig.
+3, given a reference image, WonderJourney and LucidDreamer seem to generate 3D scenes with artifacts, while our method can generate more semantic relevant to the given image, maintaining the multi-view consistency.
+In Fig.
+5, with textual prompt input, WonderJourney can inevitably generate artifacts.
+While LucidDreamer and Text2NeRF are likely to produce distorted or inconsistent viewpoint images.
 Consequently, our method demonstrates superior qualitative performance compared to these baseline approaches.
-Prompt WonderJourney[93] LucidDreamer[12] Ours
-CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓ CLIP↑ Contextual↓ FID↓
-Image 88.201 1.539 156.619 91.034 1.918 156.651 93.565 1.538 89.829
-Table 1: Quantitative comparison results of our method with the baseline WonderJourney and LucidDreamer, all capable of processing a single images or text.
-↑ means the higher, the better, ↓ means the lower, the better.
-Prompt WonderJourney[93] LucidDreamer[12] Text2NeRF[95] ProlificDreamer[86] SJC[84] Ours
-CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑ CLIP-Similarity↑
-Text 27.480 26.323 28.695 25.245 24.152 30.544
-Table 2: Quantitative comparison results of our method with the baseline WonderJourney, LucidDreamer, Text2NeRF, ProlificDreamer, and SJC, all of which only accept text inputs.
-↑ means the higher, the better, ↓ means the lower, the better.
 4.3 Ablation Studies and Analysis
@@
 To verify the effectiveness of the Consistency-Enhanced MAE, we conduct ablation studies on different strategies: removing the MAE scene completion, denoted as w/o MAE; removing the masked VQ-VQE codebook, denoted w/o codebook; removing both MAE scene completion and the masked VQ-VAE codebook, denoted as w/o MAE&\&&codebook.
+For w/o MAE, we utilize the stable diffusion (sd) inpainting to complete the scene, which is utilized in LucidDreamer.
 As shown in Tab.
@@
 6 demonstrates visually that when the model removes either of the MAE and codebook modules, it causes inconsistencies between the different views.
-Method Image Text
-CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
-w/o MAE 88.284 1.722 124.242 28.598
-w/o codebook 87.232 1.827 134.568 28.023
-w/o MAE&\&&codebook 82.392 2.024 155.322 26.767
-full model 93.565 1.538 89.829 30.544
-[…]
-tableAblation experiments regarding the MAE module and the codebook.
+This superiority of our Consistency-Enhanced MAE in handling detailed complementation is evident in Fig.
+6.
+Method CLIP↑ BRISQUE↓ NIQE↓
+w/o MAE 27.745 41.243 6.024
+w/o codebook 27.983 38.335 5.834
+w/o MAE&\&&codebook 26.674 44.234 6.653
+full model 30.544 20.324 2.522
+Table 2: Ablation experiments regarding the MAE module and the codebook.
 Figure 6: Results of the effectiveness of the MAE module and the codebook.
@@
 The full model achieves the lowest FID values, while similar in other metrics, indicating the 3D-aware generative refinement module plays an important role in providing intricate geometry and texture details.
-Method Image Text
-CLIP↑ Contextual↓ FID↓ CLIP-Similarity↑
-w/o GAN loss 92.324 1.563 156.244 30.422
-w/o video-assisted 93.533 1.548 94.294 30.242
-full model 93.565 1.538 89.829 30.544
-[…]
-tableAblation experiments regarding the 3D-aware generative refinement module.
+Method CLIP↑ BRISQUE↓ NIQE↓
+w/o GAN loss 25.234 33.234 7.234
+w/o video-assisted 28.341 24.342 5.342
+full model 30.544 20.324 2.522
+Table 3: Ablation experiments regarding the 3D-aware generative refinement module.
 Figure 7: Results of the effectiveness of the 3D-aware generative refinement module.
@@
 However, the optimization process remains an area for potential enhancement, as it is currently constrained by our use of implicit neural fields to enforce physical consistency.
-In our future research, we aim to explore faster, more easily optimized, and sustainable three-dimensional representations of scenes.
+In our future research, we aim to explore faster, more easily optimized, and sustainable 3D scene representations.
 These will serve as mechanisms to articulate consistency surface constraints, thus accelerating the generation process.
 Appendix A Appendix
-A.1 Additional Implementation Details
-A.1.1 GAN-based Training Strategy Details
-For the incorporated discriminator, we adopt a similar architecture, regularization function, and loss weight as the EG3D model [7], with some distinctions.
-The vanilla structure of a 3D GAN involves a 3D generator that incorporates a super-resolution module, accompanied by a discriminator that accepts both coarse and fine image inputs [7].
+A.1 Implementation Details
+We implement our Scene123 with the Pytorch framework (Paszke et al.
+2019) and adopt TensoRF (Chen et al.
+2022) as the radiance field.
+To ensure TensoRF can accommodate scene generation over a large view range, we position the camera near the center of the NeRF bounding box and configure it with outward-facing viewpoints.
+The dimension of the masked codebook is 2048$\times$16.
+For only text prompt input, we utilize the stable diffusion model in version 2.0 (Rombach et al.
+2022) to generated initial image $\mathbf{I_{0}}$ related to the input prompt.
+Moreover, for depth estimation, we use the boosting monocular depth estimation method (Miangoleh et al.
+2021) with pre-trained LeReS model (Yin et al.
+2021) to estimate the depth for each view.
+For the image-to-video pipeline, we opt for the Stable Video Diffusion in version SVD-XT, which is the same architecture as SVD (Blattmann et al.
+2023a) but finetuned for 25 frame generation.
+During training, we use the same setting as (Chen et al.
+2022) for the optimizer and learning rate and set the hyperparameters in our objective function as $\lambda_{d}=0.005$, $\lambda_{t}=0.001$, $\lambda_{dist}=0.001$.
+GAN-based Training Strategy Details.
+For the incorporated discriminator, we adopt a similar architecture, regularization function, and loss weight as the EG3D model (Chan et al.
+2022), with some distinctions.
+The vanilla structure of a 3D GAN involves a 3D generator that incorporates a super-resolution module, accompanied by a discriminator that accepts both coarse and fine image inputs (Chan et al.
+2022).
 Notably, due to the contextual disparities between text-to-3D and 3D GAN applications, we opt to omit the super-resolution component from the 3D GAN architecture.
 This choice stems from the consistent need to extract mesh or voxel representations from the 3D model within the context of text-to-3D.
-A.2 Additional Qualitative Results
-We provide additional qualitative results in Fig.
-8, including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
-Figure 8: More style visualization examples generated by Scene123.
-We provide additional results including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes, and fairy-tale scenes with realistic renderings and precise depth details.
+The camera motion in the SVD model is sometimes limited (Blattmann et al.
+2023a), making it unsuitable for directly training the NeRF model.
+Instead, we use our 3D-Aware Generative Refinement approach, which leverages a discriminator to optimize the process.
+In this setup, the video diffusion model’s output is treated as real data, while the 3D neural radiance field model’s renderings are treated as fake data.
+Despite the small range of camera motion in the images generated by SVD, the adversarial training process provides necessary regularization.
+The discriminator learns to distinguish between real support set images and fake renderings, thereby pushing the neural radiance field model to produce outputs that are indistinguishable from real images.
+This continuous adversarial interaction compels the generator to improve by learning finer details and more accurate textures, ultimately enhancing the quality of the 3D scene generation.
+This method ensures that even with limited camera motion, the GAN loss remains effective and contributes to generating high-quality 3D scenes, as shown in Fig.
+7 in the manuscript.
+Depth alignment details.
+We use the depth estimation model $f_{e}$ to estimate the depth map $D_{k}^{E}$ for the initial view $\mathbf{I_{0}}$.
+Note that, unlike the depth map $D_{0}$ of the initial view, $D_{k}^{E}$ cannot be directly taken as the supervision to update the radiance field since it is predicted independently and could conflict with known depth maps such as $D_{k}^{R}$ in the overlapping regions.
+To solve this issue, we implement depth alignment to align the estimated depth map to the known depth values in the radiance field (Zhang et al.
+2024b).
+Due to the lack of geometric constraint during the depth estimation, the predicted depth values could be misaligned in the overlapping regions (Luo et al.
+2020), for example, the estimated depth $D_{k}^{E}$ of the inpainted view may be inconsistent with the depth $D_{k}^{R}$ rendered from NeRF since $D_{k}^{R}$ is constrained by previous known views.
+The inconsistency is manifested in two aspects: scale difference and value difference.
+For instance, the distance difference of two pixel-aligned spatial points and the depth value of a specific point could be both different in depth maps estimated from different views.
+The former is the scale difference and the latter is the value difference.
+In the case of scale difference, we cannot align both points by shift processing because even if we align the depth value of one of the points, the other point is still misaligned.
+To eliminate the scale and value differences between the overlapping regions of the rendered depth map $D_{k}^{R}$ and the estimated depth map $D_{k}^{E}$ of the novel view, we introduce a two-stage depth alignment strategy.
+Specifically, we first globally align these two depth maps by compensating for mean scale and value differences.
+Then we finetune a pre-trained depth alignment network to produce a locally aligned depth map.
+To determine the mean scale and value differences, we first randomly select $M$ pixel pairs from the overlapping regions and deduce their 3D positions under depth $D_{k}^{R}$ and $D_{k}^{E}$, denoted as $\left\{(\mathbf{x}_{j}^{R},\mathbf{x}_{j}^{E})\right\}_{j=1}^{M}$.
+Next, we calculate the average scaling score $s$ and depth offset $\delta$ to approximate the mean scale and value differences:
+$s=\frac{1}{M-1}\sum_{j=1}^{M-1}\frac{\|\mathbf{x}_{j}^{R}-\mathbf{x}_{j+1}^{R}\|_{2}}{\|\mathbf{x}_{j}^{E}-\mathbf{x}_{j+1}^{E}\|_{2}},$ (6)
+$\delta=\frac{1}{M}\sum_{j=1}^{M}\left(z\left(\mathbf{x}_{j}^{R}\right)-z\left(\mathbf{\hat{x}}_{j}^{E}\right)\right),$ (7)
+where $\mathbf{\hat{x}}_{j}^{E}=s\cdot\mathbf{x}_{j}^{E}$ indicates the scaled point and $z(\mathbf{x})$ represents the depth value of point $\mathbf{x}$.
+Then $D_{k}^{E}$ can be globally aligned with $D_{k}^{R}$ by $D_{k}^{global}=s\cdot D_{k}^{E}+\delta$.
+Since depth maps used in our pipeline are predicted by a network, the differences between $D_{k}^{R}$ and $D_{k}^{E}$ are not linear, that is why the global depth aligning process cannot solve the misalignment problem.
+To further mitigate the local difference between $D_{k}^{global}$ and $D_{k}^{R}$, we train a pixel-to-pixel network $f_{\psi}$ for nonlinear depth alignment.
+During optimization of each view, we optimize the parameter $\psi$ of the pre-trained depth alignment network $f_{\psi}$ by minimizing their least square error in the overlapping regions:
+$\mathop{\min}_{\psi}\left\|\left(f_{\psi}(D_{k}^{global})-D_{k}^{R}\right)\odot M_{k}\right\|_{2}.$ (8)
+Finally, we can derive the locally aligned depth using the optimized depth alignment network: $\hat{D}_{k}=f_{\psi}(D_{k}^{global})$.
+Figure 8: Comparisons of SD inpainting and our Consistency-Enhanced MAE in handling detailed complementation.
+A.2 Additional analysis
+Comparison between SD inpainting and Our Consistency-Enhanced MAE in detail completion.
+Our Consistency-Enhanced MAE ensures consistency across views by using a learnable codebook to distribute global information to every invisible area.
+Traditional VAE methods quantize the latent space, leading to artifacts at the boundaries between quantized latents.
+In contrast, the cross attention mechnism allows for dynamic and flexible feature weighting, better captures global context, and preserves detailed information without the loss associated with quantization.
+The learnable codebook continuously adapts during training, enhancing the model’s stability and performance in generating high-quality 3D scenes.
+However, the stable diffusion (SD) model, which is originally utilized to inpaint images in LucidDreamer (Chung et al.
+2023), is less effective for the multi-view complementation task.
+While SD can achieve complementation through conditional diffusion, it struggles with complex structures and view-consistency tasks, as it is not tailored for this purpose.
+Although SD performs well in generating new images, it may not be as effective in complementing local details compared to MAE, which is specifically designed for this task.
+This superiority of our Consistency-Enhanced MAE in handling detailed complementation is evident in Fig.
+8.
+Video Generation Model promotes our Scene123’s quality.
+The 3D scenes generated through our 3D-Aware Generative Refinement module are of higher quality than the image-to-video generation pipeline.
+As depicted in Fig.
+9, our Scene123 does not strictly require a high-quality support set.
+The novel views generated by our method are generally superior to those produced by the image-to-video pipeline.
+This is because our 3D-Aware Generative Refinement uses a discriminator to manage optimization, distinguishing between real and fake images.
+This adversarial interaction improves the generator’s ability to learn finer details and more accurate textures, enhancing the overall quality of 3D scene generation.
+This demonstrates our approach’s adaptability in handling image-to-video generation failures.
 A.3 User Study
-For completeness, we follow previous works [86, 93] and conduct a user study by comparing Scene123 with WonderJourney [93], LucidDreamer [12] under 40 image prompts, 20 image prompts for each baseline.
-For text prompt input, we compare Scene123 with Text2NeRF [95], ProlificDreamer [86], SJC [84] under 60 text prompts, 20 text prompts for each baseline.
-The volunteers are shown the generated results of our Scene123 and baselines and asked to choose the better one in terms of fidelity, details and vividness.
-We collect results from 39 volunteers, yielding 3120 pairwise comparisons.
+For completeness, we follow previous works (Wang et al.
+2024; Yu et al.
+2023) and conduct a user study by comparing Scene123 with WonderJourney (Yu et al.
+2023), LucidDreamer (Chung et al.
+2023) under 40 image prompts, 20 image prompts for each baseline.
+For text prompt input, we also compare Scene123 with WonderJourney (Yu et al.
+2023), LucidDreamer (Chung et al.
+2023) and Text2NeRF (Zhang et al.
+2024b) under 60 text prompts, 20 text prompts for each baseline.
+The participants are shown the generated results of our Scene123 and baselines and asked to choose the better one in terms of fidelity, details and vividness.
+We collect results from 39 participants, yielding 3120 pairwise comparisons.
 The results are shown in Tab.
-3.
+4.
 Our method outperforms all of the baselines.
-Method WonderJourney LucidDreamer Text2NeRF ProlicDreamer SJC
-Prefer baseline 31.12 23.42 46.27 10.25 5.50
-Prefer Scene123 (Ours) 68.88 76.58 53.73 89.75 94.50
+Figure 9: Comparisons between support set generated by image-to-video pipeline and the generated views of Our Scene123.
+Method WonderJourney LucidDreamer Text2NeRF
+Prefer baseline 31.12 23.42 46.27
+Prefer Scene123 (Ours) 68.88 76.58 53.73
 Table 4: Results of user study. The percentage of user preference (↑) is reported in the table.
-A.4 Additional analysis
-As depicted in Fig.
-9, Scene123 doesn’t strictly require a high-quality support set.
-The novel views generated by our method are generally superior to the views generated by the image-to-video pipeline.
-This illustrates the adaptability of our approach to cope with such instances of generation failure.
-Figure 9: Comparisons between support set generated by image-to-video pipeline and the generated views of Our method.
+A.4 More Qualitative Results
+We provide more qualitative results in Fig.
+10, Fig.
+11, Fig.
+12, Fig.
+13, including indoor scenes, outdoor scenes, outdoor buildings, object-centered scenes with realistic renderings and precise depth details.
+Part 1 / 4
+Figure 10: More qualitative examples generated by our Scene123 from a single image input.
+Part 2 / 4
+Figure 11: More qualitative examples generated by our Scene123 from a single image input.
+Part 3 / 4
+Figure 12: More qualitative examples generated by our Scene123 from a text prompt input.
+Part 4 / 4
+Figure 13: More qualitative examples generated by our Scene123 from a text prompt input.